The value V of a motor cycle that is t years old can be modeled by the function
V = 16,500(0.82)^t
. According to the model, when will the motor cycle be
worth $4000?

Answer by Boreal(15235)

4000=16500(0.82^t)
divide both sides by 16500
0.2424=0.82^t
ln both sides
-1.4171=t ln 0.82
t= -1.4171/ln 0.82
=(if I round only at the end)
t=7.14 years